OpenStudy (anonymous):

There were 150 freshman and sophomores at the high school musical Reese and her friends attended. Freshman tickets were $2.00 each and sophomore tickets were $3.00 each. The receipts for freshman and sophomore tickets totaled $400. How many sophomores bought tickets? Answer 75 100 25 50

OpenStudy (mathmate):

Do this in your head: If only freshmen attended, there would be only $300. The extra $100 is generated by replacing 100 freshmen tickets with sophomore tickets. So there are 100 sophomores and 50 freshmen.

OpenStudy (anonymous):

We know that the total amount of people present is 150 (including sophmores and freshmen) So, let # of sophmores attending = S # of freshmen attending = F So:1) S + F = 150 We also know that for each Freshman, the ticket costs $2.00 each And for Sophmores, they cost $3.00 per person And we know that in total, $400 was spent. So 2) 2F + 3S = 400 Rearranging 1) for S gives us: S = 150-F And substitution that into 2), gives us 2F + 3(150 - F) = 400 Isolating for F, we get: F = 50, and using that number in 1) S = 100